OpenStudy (yooooooo):

the slope of the curve y^2-xy-3x=1 at the point (0 -1) is

OpenStudy (anonymous):

2y (dy/dx)-y-x(dy/dx)-3=0 dy/dx=(y+3)/(2y-x) at 0,-1 slope will be=-1